OpenStudy (anonymous):

Which lists all the integer solutions of the equation |x| = 4? A. –4 and 4 B. 0 and 4 C. –4 only D. 4 only

OpenStudy (michele_laino):

hint: by definition of absolute value of \( \large x \), we have: \[\Large \left| x \right| = \left\{ {\begin{array}{*{20}{c}} {x,\quad x \geqslant 0} \\ { - x,\quad x < 0} \end{array}} \right.\]

OpenStudy (anonymous):

i think A